OpenStudy (anonymous):

what is improper fraction of 27/9

OpenStudy (triciaal):

is that whole number 2 with 7/9?

OpenStudy (triciaal):

2 + 7/9 =2/1 +7/9 lcd = 9 (18 + 7)/9 = 25/9 note 27/9 is already improper =3